Grade 11Physical Chemistry

What volume of 0.4 M FeCl3.6H2O will contain 600mg of Fe3+ ?

devika azhakesan
10 Years agoGrade 11

1 Answer

Dhaval
8 Years ago
Chemistry help needed! What Volume of 0.4 M Fecl3.6H20 will contain 600 mg of Fe3+?2 ANSWERSKumaraswami SathiavasanKumaraswami Sathiavasan, MSc in Chemistry & IAS officer(retd.)Answered Sep 19, 2015Molar mass of FeCl3.6H2O = 55.85 + 35.45 x 3 + 18 x 6 = 270.2 gmAmount of the salt in 1 litre of 1M solution = 270.2 gmAmount of the salt in 1 litre of 0.4M solution = 270.2 x 0.4 =1o8.1 gmAtomic mass of iron (Fe) = 55.85 u; Molar mass of iron = 55.85 gm.One mole of FeCl3.6H2O contains only one mole of Fe3+ (iron). That is, 270.2 gm of the salt contains 55.85 gm of iron (Fe3+).Amount of Fe3+ (iron) in 1 litre of 0.4M solution = 108.1 x 55.85/270.2 = 22.34 gm. So, 1 litre (1000 ml) of the 0.4M salt solution contains 22.34 gm of Fe3+ (iron).Volume needed for 600 mg (0.6 gm) of Fe3+ = 0.6 x 1000/22.34 = 26.9 ml.